# EFEAR-4D: Ego-Velocity Filtering for Efficient and Accurate 4D radar Odometry

### v1

@@ Abstract
Odometry is a crucial component for successfully implementing autonomous navigation, relying on sensors such as cameras, LiDARs and IMUs. However, these sensors may encounter challenges in extreme weather conditions, such as snowfall and fog. The emergence of FMCW radar technology offers the potential for robust perception in adverse conditions. As the latest generation of FWCW radars, the 4D mmWave radar provides point cloud with range, azimuth, elevation, and Doppler velocity information, despite inherent sparsity and noises in the point cloud. In this paper, we propose EFEAR-4D, an accurate, highly efficient, and learning-free method for large-scale 4D radar odometry estimation. EFEAR-4D exploits Doppler velocity information delicately for robust ego-velocity estimation, resulting in a highly accurate prior guess. EFEAR-4D maintains robustness against point-cloud sparsity and noises across diverse environments through dynamic object removal and effective region-wise feature extraction. Extensive experiments on two publicly available 4D radar datasets demonstrate state-of-the-art reliability and localization accuracy of EFEAR-4D under various conditions. Furthermore, we have collected a dataset following the same route but varying installation heights of the 4D radar, emphasizing the significant impact of radar height on point cloud quality - a crucial consideration for real-world deployments. Our algorithm and dataset will be available soon at https://github.com/CLASS-Lab/EFEAR-4D.
@@
4D mmWave radar, radar odometry, ego-velocity estimation
@@ I Introduction
Figure 1: Our efficient 4D radar odometry and mapping.
@@
To tackle these challenges, this paper introduces a novel method for 4D radar odometry estimation, EFEAR-4D (Ego-velocity Filtering for Efficient and Accurate 4D radar Odometry), as shown in Fig. 1. Specifically, we associate direction with Doppler velocity for each point and use a density-based clustering method DBSCAN [11] to filter out noises and remove point cloud of dynamic objects, followed by ego-velocity estimation leveraging the geometric distribution characteristics of Doppler velocity from all static objects. Considering sparsity in radar point cloud, we conduct region-wise geometric and intensity-based feature extraction for registration. And Point-to-Point(P2P) registration is employed for computing odometry. Moreover, To investigate how limited FOV of 4D radars affects the numbers of point cloud and thus the performance of odometry estimation, we design experiments with 4D radar positioned at various heights, and provide comparative analysis as a guidance for future 4D radar real-world deployments. Our main contributions are summarized as follows:
@@
    The proposed odometry method CFEAR-4D is thoroughly validated across various scales, vehicles, and environments, consistently demonstrating high performance in diverse scenarios. A dataset with 4D radar installation at different heights is complied. The dataset and code are open-sourced to facilitate related research in the community.
@@ II Related Works
We will briefly summarize related works from two aspects: traditional radar (2D or 3D) and 4D imaging radar odometry.
@@ II-A Traditional radar odometry
Radar-based odometry methods mostly can be divided into two categories: feature-based methods and direct methods. Feature-based methods typically extract and match distinctive features for registration between consecutive scans. Cen et al. [12] recognize landmarks as features to perform data association. In radarSLAM [6], SURF descriptor is extracted for pose estimation. Adolfsson et al. [1] incorporate normal vector and other factors as weights for point-to-line (P2L) scan matching. For direct methods, [13] uses the Fourier-Mellin transform to recover all rigid transformation parameters. ICP and NDT are used for radar scan registration in [14] and [15], respectively.
@@
The Doppler velocity information of 4D radars provides a new way to estimate ego velocity that traditional 2D radars cannot. With accurate ego-velocity estimation as a reliable motion prior, pose estimation can be enhanced significantly. The simple idea of ego-velocity estimation is to consider the Doppler velocity of each stationary point as a projection of the true ego-velocity in the radial direction between the stationary point and the center of the radar. Consequently, the least squares methods can be used to calculate the ego-velocity. However, substantial noises emerge in 4D radar velocity measurements, particularly in scenarios involving dynamic objects, where the obtained Doppler velocity represents the relative velocity between two moving entities.
Kramer et al. [16] introduce Cauchy robust norm to address this challenge. Christopher and Gert [17] utilize RANSAC method to mitigate the effects of outliers. Though it works well with a small number of points, it performs poorly in large-scale scenarios, where radar point cloud typically with thousands of points. Chen et al. [8] present a 2D ego-velocity estimation method by refining ground points, however, such a method does not work in 4D cases. Due to sparsity in 4D radar point cloud, many frames lack sufficient ground points. Zeng et.al. [18] present a signal model for radar ego-motion estimation in the presence of velocity ambiguity. However, it concentrates on the fuzziness of the radar point cloud itself rather than the impact of factors such as dynamic objects in real-world situations. Therefore, we aim at proposing a novel ego-velocity filtering method, which provides accurate ego-velocity estimates by filtering out significant noises in diver dynamic scenarios.
@@ III-A Ego-velocity filtering
Doppler velocity represents the relative velocity on the radial direction between the measured point and the center of the radar. Therefore, if the measured point is a stationary point, the Doppler velocity can be mathematically expressed as:
@@
where $\textbf{v}^{R}$ and $\textbf{v}^{S}$ denote the vectors of Doppler velocity and ego-velocity, $|\textbf{v}^{R}|$ denotes the value of vector $\textbf{v}^{R}$, and $r$ is the unit vector which indicates the direction of the Doppler velocity. In the default radar body frame, the $x$ and $y$ coordinates can be expressed as below:
$V_{x}$ $=|\textbf{v}^{R}|\sin\theta=\frac{|\textbf{v}^{S}|}{2}\sin{2\theta}$ (2)
$V_{y}$ $=|\textbf{v}^{R}|\cos\theta=\frac{\cos{2\theta}+1}{2}|\textbf{v}^{S}|$ (3)
As a result, we can obtain that
$V^{2}_{x}+\Big{(}V_{y}-\frac{|\textbf{v}^{S}|}{2}\Big{)}^{2}=\Big{(}\frac{|\textbf{v}^{S}|}{2}\Big{)}^{2}$ (4)
From a geometric perspective, the above equation (4) indicates that the Doppler velocity distribution for stationary points forms a circle in the XY plane and $|\textbf{v}^{S}|$ represents the diameter through the origin, as illustrated in Fig. 3. Every point on the circle can be regarded as the projection of $\textbf{v}^{S}$, reflected from a stationary place. Thus, fitting it allows us to determine the real ego velocity. In the ground vehicle setting, given that the velocity along the Z-axis is approximately zero, we directly fit a circle in the the XY plane instead of a sphere in the 3-dimensional Euclidean space. We then employ the least squares method to compute the radius $\frac{|\textbf{v}^{S}|}{2}$. The minimization function is
$\mathop{\arg\min}\limits_{r}\sum_{i}(v^{R_{x}}_{i}-0)^{2}+(v^{R_{y}}_{i}-r)^{2}-r^{2}$ (5)
where $r=\frac{|\textbf{v}^{S}|}{2}$.
@@
However, due to the presence of dynamic objects and noises in radar point cloud, applying the least squares method directly to all data cannot yield correct ego velocity estimates. As illustrated in Fig. 3(d), the light green points correspond to the white car driving in the opposite direction of the ego vehicle, and the dark points correspond to the pedestrians, the black car in front and some other moving objects shown in Fig. 3(a). To differentiate between dynamic points, noise, and stationary points, we employ DBSCAN [11], a density-based clustering method that utilizes region growing. Although noise and dynamic objects may not be entirely distinguishable, stationary points will dominate, allowing for the straightforward identification of stationary points and facilitating more accurate estimation. Dynamic points are then removed according to the classification results of DBSCAN for better registration in the subsequent steps.
@@ III-B Region-wise feature extraction
Since point cloud of 4D radar is noisy, direct registration with all the points would lead to large error, especially in the Z-axis. In this section, we discuss how to extract some features for registration. Beginning with the filtered points, our initial step involves partitioning them into voxels. And for the $i$-th voxel $V_{i}$, we utilize KD-tree [22] to search for its neighboring points, which include not only points within this voxel but also adjacent points outside the area, denoted as $P_{i}$. Subsequently, employing intensity weights, we calculate the weighted mean point for $P_{i}$ as the feature point of the voxel. The weighted sample covariance matrix is denoted as $C_{i}$.
Besides, since the intensity information of 4D radar is relatively reliable, if the maximum intensity within $P_{i}$, represented as $w_{i}$, exceeds a threshold $w_{thre}$, we select the corresponding point as a feature point, which we empirically find alleviates the problem of Z-axis drift for 4D radar to a certain extent. The final feature point for $P_{i}$ is denoted as $p_{i}$. Moreover, to improve noise immunity, we perform principal component analysis (PCA) [23] to compute normal vector $u_{i}$ as another feature of the voxel $V_{i}$. The eigenvalues and eigenvectors of covariance matrix $C$ are calculated as follows:
@@
To the best of our knowledge, it is the first time to use intensity information to extract features in 4D radar odometry. Furthermore, choosing a feature point and normal vector to represent a voxel lightens the influence of noise for registration as noise points in 4D radar point cloud are typically unstable and with low intensity.
@@
Ours 2.4282 3.2039 0.8566 15.5033 14.3756 8.4184 17.9600 1.1431
Figure 4: The current scan(a) and corresponding keyframe (b) are projected in XY plane. red points in (a) and black points in (b) are successfully matched feature points in two scans, respectively. The real current point cloud and extracted normal vectors are shown in (c). For visualization, the size of voxels here is 10m, with features being much denser in quantity during actual operation.
@@ III-C Scan-to-map matching
Leveraging the ego-velocity computed in the preceding step, we derive the initial guess of pose ${}^{w}\hat{T_{k}}$, and align it into sub-map coordinates for the current frame, denoted as ${}^{s}\hat{T_{k}}$. Each voxel in the current frame searches for its corresponding voxel in the sub-map in a certain radius. For corresponding voxels, feature points are then matched and a filtering process is applied to exclude points with a normal vector dot product $v_{i}\cdot v_{j}<z_{thre}$, indicating potential false association. As a result, all matching points ($\prescript{s}{}{p},\prescript{k}{}{p}$) form a sparse set $\Omega$ as shown in Fig. 4(a) and (b).
@@ IV-A1 Comparative evaluation with MSC dataset
(a) RURAL_A2
(b) RURAL_D2
(c) URBAN_A0
(d) URBAN_B0
(e) URBAN_D0
(f) URBAN_F0
@@
We present the odometry error in Tab. I and depict the trajectory results in Fig. 6. Unexpectedly, we observe that our errors are lower than those of ORB-SLAM3 across nearly all scenarios and are approaching the level of LiDAR, or even surpassing it in some scenarios. In static environments (U_A0, U_B0), owing to precise ego velocity estimation, our EFEAR-4D method converge to a relatively accurate pose estimate, outperforming other comparative methods. In sequence U_B0, the dense vegetation makes it challenging for LiDAR to distinguish columnar features, leading to degradation in the performance of LEGO-LOAM. The other two radar methods exhibit significant errors due to the lack of feature extraction. In high-speed environment, rapid motion causes a large distance between the point cloud of consecutive frames, resulting in fewer features available for matching. As the FOV of LiDAR is much larger than radar, LEGO-LOAM demonstrates superior stability and performance. Similar to radar, the limited FOV of camera leads to rapid performance degradation of ORB-SLAM3.
In dynamic environments, by removing dynamic objects and benefiting from the weak reflection on pedestrians or other dynamic objects, our method is minimally affected by dynamic objects, unlike other methods, particularly vision-based and LiDAR-based approaches. Furthermore, compared with GICP and APDGICP, our method extracts features and filters unstable points based on intensity, further enhancing our odometry performance. In snowy environments, particularly in R_D2, where LEGO-LOAM occasionally fails to register frames, radar-based methods outperform those utilizing other sensors, demonstrating the efficacy of radar in all weather conditions.
Fig. 7 describes the relative error of ORB-SLAM3, LEGO-LOAM, baseline radar odometry, and our radar odometry EFEAR-4D using different path lengths and speeds with different sequences, following the popular KITTI odometry evaluation protocol. Our EFEAR-4D shows a stable performance on all the datasets except in high-speed scenarios where it is slightly inferior to LEGO-LOAM. Both 4DRadarSLAM (apdgicp) and ORB-SLAM3 have a relatively large error and their performances are sensitive to speed variations. It is worth pointing out that both our EFEAR-4D and 4DRadarSLAM perform well in dynamic environments, proving that 4D radar optometry is not less sensitive to moving objects than its counterparts of vision and LiDar.

### v2

@@ Abstract
Odometry is a crucial component for successfully implementing autonomous navigation, relying on sensors such as cameras, LiDARs and IMUs. However, these sensors may encounter challenges in extreme weather conditions, such as snowfall and fog. The emergence of FMCW radar technology offers the potential for robust perception in adverse conditions. As the latest generation of FWCW radars, the 4D mmWave radar provides point cloud with range, azimuth, elevation, and Doppler velocity information, despite inherent sparsity and noises in the point cloud.  EFEAR-4D exploits Doppler velocity information delicately for robust ego-velocity estimation, resulting in a highly accurate prior guess. EFEAR-4D maintains robustness against point-cloud sparsity and noises across diverse environments through dynamic object removal and effective region-wise feature extraction. Extensive experiments on two publicly available 4D radar datasets demonstrate state-of-the-art reliability and localization accuracy of EFEAR-4D under various conditions. Furthermore, we have collected a dataset following the same route but varying installation heights of the 4D radar, emphasizing the significant impact of radar height on point cloud quality, a crucial consideration for real-world deployments. Our algorithm and dataset are available at https://github.com/CLASS-Lab/EFEAR-4D.
@@
SLAM, Autonomous Vehicle Navigation, 4D mmWave radar, Radar Odometry
@@ I Introduction
Figure 1: Our efficient 4D radar odometry and mapping. We test our system on our AGV platform. The mapping result and corresponding ground-truth are depicted on the right.
@@
To tackle these challenges, this paper introduces a novel method for 4D radar odometry estimation, EFEAR-4D (Ego-velocity Filtering for Efficient and Accurate 4D Radar-odometry), as shown in Fig. 1. Specifically, we associate direction with Doppler velocity for each point and use a density-based clustering method DBSCAN [11] to filter out noises and remove point cloud of dynamic objects, followed by ego-velocity estimation leveraging the geometric distribution characteristics of Doppler velocity from all static objects. Considering sparsity in radar point cloud, we conduct region-wise geometric feature extraction for registration. And Point-to-Point(P2P) registration is employed for computing odometry. Moreover, To investigate how limited FOV of 4D radars affects the numbers of point cloud and thus the performance of odometry estimation, we design experiments with 4D radar positioned at various heights, and provide comparative analysis as a guidance for future 4D radar real-world deployments. Our main contributions are summarized as follows:
@@
    The proposed odometry method EFEAR-4D is thoroughly validated across various scales, vehicles, and environments, consistently demonstrating high performance in diverse scenarios. A dataset with 4D radar installation at different heights is complied. The dataset and code are open-sourced to facilitate related research in the community.
@@ II Related Works
We will briefly summarize related works from two aspects: traditional radar (2D, including the $x$ and $y$ axes, or 3D, including the $x,y$ axes, and Doppler velocity) and 4D imaging radar (including the $x,y,z$ axes and Doppler velocity) odometry.
@@ II-A Traditional radar odometry
Radar-based odometry methods mostly can be divided into two categories: feature-based and direct methods. Feature-based methods typically extract and match distinctive features for registration between consecutive scans. Cen et al. [12] recognize landmarks as features to perform data association. In radarSLAM [6], SURF descriptor is extracted for pose estimation. Adolfsson et al. [1] incorporate normal vector and other factors as weights for point-to-line scan matching. For direct methods, [13] uses the Fourier-Mellin transform to recover all rigid transformation parameters. ICP and NDT are used for radar scan registration in [14] and [15], respectively.
@@
The Doppler velocity information from 3D/4D radars offers a novel method to estimate ego velocity, which traditional 2D radars cannot achieve. Recently, several works have proposed Doppler velocity estimation algorithms for spinning radars [16] [17]. However, the extension of these methods to ego velocity estimation remains an open research question. With accurate ego-velocity estimation as a reliable motion prior, pose estimation can be enhanced significantly. The simple idea of ego-velocity estimation is to consider the Doppler velocity of each stationary point as a projection of the true ego-velocity in the radial direction between the stationary point and the center of the radar. Consequently, the least squares methods can be used to calculate the ego-velocity. However, substantial noises emerge in 4D radar velocity measurements, particularly in scenarios involving dynamic objects, where the obtained Doppler velocity represents the relative velocity between two moving entities.
Kramer et al. [18] introduce Cauchy robust norm to address this challenge. Christopher and Gert [19] utilize RANSAC method to mitigate the effects of outliers. Though it works well with both small and large numbers of points [20], its performance deteriorates in complex environments with crowded dynamic objects, due to the lack of identifying and removal of dynamic objects in the point cloud. Chen et al. [8] present a 2D ego-velocity estimation method by refining ground points, however, such a method does not work in 4D cases. Due to sparsity in 4D radar point cloud, many frames lack sufficient ground points. Zeng et.al. [21] present a signal model for radar ego-motion estimation in the presence of velocity ambiguity. However, it concentrates on the fuzziness of the radar point cloud itself rather than the impact of factors such as dynamic objects in real-world situations. Therefore, we aim at proposing a novel ego-velocity filtering method, which provides accurate ego-velocity estimates by filtering out significant noises in diverse dynamic scenarios.
@@ III-A Ego-velocity filtering
Doppler velocity is the relative velocity in the radial direction from the center of the radar to the measured point. Given that the measured point is a stationary point, we have:
@@
where $\textbf{v}^{R}\in\mathbb{R}^{3}$ is the Dopper velocity (in the radial direction from the center of the radar to the measured point) and $\textbf{v}^{S}\in\mathbb{R}^{3}$ is the ego-velocity vector, $\textbf{r}\in\mathbb{R}^{3}$ is a unit vector along the direction of the observed point in the radar coordinate system, indicating the direction of the Doppler velocity, and $\theta$ is the angle from $\textbf{v}^{R}$ to $\textbf{v}^{S}$, ranging from $-\pi$ to $\pi$.
Given that the velocity along the z-axis is negligible in ground vehicle settings, we simplify the model by directly setting the z-axis velocity to zero. As depicted in Fig. 3 (c), in the default radar body frame $V_{x}$-$V_{y}$-$O$, denote the angle from the x-axis to $\textbf{v}^{R}$ as $\beta$. and the decomposing component of $\textbf{v}^{R}$ in the $V_{x}$ and $V_{y}$ coordinates as $v_{x},v_{y}$, respectively. It follows from (1) that:
$v_{x}$ $=|\textbf{v}^{R}|\cos\beta=\frac{|\textbf{v}^{S}|}{2}(\cos{(\theta+\beta)}+\cos{(\beta-\theta)})$
$v_{y}$ $=|\textbf{v}^{R}|\sin\beta=\frac{|\textbf{v}^{S}|}{2}(\sin{(\theta+\beta)}+\sin{(\beta-\theta)})$
Defining $\gamma\triangleq\beta+\theta$, then it follows that  [^1] Noting that $\textbf{v}^{S}$ is on the circle, the other candidate equation $\Big{(}v_{x}-\frac{|\textbf{v}^{S}|}{2}\cos{(\beta-\theta)}\Big{)}^{2}+\Big{(}v_{y}-\frac{|\textbf{v}^{S}|}{2}\sin{(\beta-\theta)}\Big{)}^{2}=\Big{(}\frac{|\textbf{v}^{S}|}{2}\Big{)}^{2}$ is excluded.:
$\Big{(}v_{x}-\frac{|\textbf{v}^{S}|}{2}\cos{\gamma}\Big{)}^{2}+\Big{(}v_{y}-\frac{|\textbf{v}^{S}|}{2}\sin{\gamma}\Big{)}^{2}=\Big{(}\frac{|\textbf{v}^{S}|}{2}\Big{)}^{2}$ (2)
From a geometric perspective, the above equation (2) indicates that all 2D points $(v_{x},v_{y})$ lie on the same circle with $(\frac{|\textbf{v}^{S}|}{2}\cos{\gamma},\frac{|\textbf{v}^{S}|}{2}\sin{\gamma})$ as the center of circle and $|\textbf{v}^{S}|$ as the diameter, as illustrated in Fig. 3 (d). Thus $\textbf{v}^{S}=\begin{bmatrix}|\textbf{v}^{S}|\cos{\gamma},|\textbf{v}^{S}|\sin{\gamma},0\end{bmatrix}^{\top}$. Referring to (2), we can employ the least squares method to compute $\textbf{v}^{S}$ with the minimization function:
$\mathop{\arg\min}\limits_{x_{0},y_{0}}\sum_{i}[(v^{i}_{x}-x_{0})^{2}+(v^{i}_{y}-y_{0})^{2}-(x_{0}^{2}+y_{0}^{2})]^{2}$ (3)
where $v^{i}_{x}$ and $v^{i}_{y}$ are the $x,y$ components of Doppler velocity from the $i$-th point in the radar point cloud.
After obtaining optimal solution $x^{*}_{0}$ and $y^{*}_{0}$ from (3), it follows that:
$\textbf{v}^{S}=\begin{bmatrix}|\textbf{v}^{S}|\cos{\gamma},|\textbf{v}^{S}|\sin{\gamma},0\end{bmatrix}^{\top}=\begin{bmatrix}2x^{*}_{0},2y^{*}_{0},0\end{bmatrix}^{\top}.$
@@
However, due to the presence of dynamic objects and noises in radar point cloud, applying the least squares method directly to all data cannot yield correct ego velocity estimates. As illustrated in Fig. 3(d), the light green points correspond to the white car driving in the opposite direction of the ego vehicle, and the dark points correspond to the pedestrians, the black car in front and some other moving objects shown in Fig. 3(a). To differentiate between dynamic points, noise, and stationary points, we employ DBSCAN [11], a density-based clustering method that utilizes region growing on all pairs of $(v_{x},v_{y})$, the distribution of Doppler velocity in the horizonal plane. Although noise and dynamic objects may not be entirely distinguishable, stationary points will dominate, allowing for the straightforward identification of stationary points and facilitating more accurate estimation. Dynamic points are then removed according to the classification results of DBSCAN for better registration in the subsequent steps.
@@ III-B Region-wise feature extraction
Since the point cloud of 4D radar is noisy, direct registration with all the points would lead to large errors, especially in the $z$-axis. As such, we adapt the novel 2D radar odometry method CFEAR [1] to the characteristics of 4D radar. Following the procedure of CFEAR, the initial step is to partition the filtered points into voxels. For each voxel $V_{i}$, then a KD-tree [25] is used to search for neighbouring points, including points both within and adjacent to the voxel, denoted as $P_{i}$. Subsequently, the weighted mean point for $P_{i}$ is calculated using intensity weights, which serves as the feature point of the voxel. The weighted sample covariance matrix is denoted as $C_{i}$ and the feature point for $P_{i}$ is denoted as $p_{i}$.
Moreover, to improve noise immunity, we perform principal component analysis (PCA) [26] to compute normal vector $u_{i}$ as another feature of the voxel $V_{i}$. The eigenvalues and eigenvectors of covariance matrix $C$ are calculated as follows:
@@
Ours 2.4282 3.2039 0.8566 5.0714 8.2494 5.9241 16.4847 1.1431
Figure 4: The current scan (a) and corresponding keyframe (b) are projected in the XY plane. Red points in (a) and black points in (b) are successfully matched feature points in the two scans, respectively. The real current point cloud and extracted normal vectors are shown in (c). For visualization, the size of voxels here is 10m, with features being much denser in quantity during actual operation.
@@ III-C Scan-to-map matching
Utilizing the ego-velocity calculated in the previous step, we derive the initial pose estimate ${}^{w}\hat{T_{k}}$ based on a constant velocity model and align it with the sub-map coordinates for the current frame, denoted as ${}^{s}\hat{T_{k}}$. Therefore, if the velocity estimation is accurate, the coordinates of the same reflective points in the sub-map and the current frame should be close. Then each voxel in the current frame can search for its corresponding voxel in the sub-map in a certain radius. For corresponding voxels, feature points are then matched and a filtering process is applied to exclude points with a normal vector dot product $u_{i}\cdot u_{j}<z_{thre}$, indicating potential false association. As a result, all matching points ($\prescript{s}{}{p},\prescript{k}{}{p}$) form a sparse set $\Omega$ as shown in Fig. 4 (a) and (b).
@@ IV-A1 Comparative evaluation with MSC dataset
(a) URBAN_B0
(b) URBAN_F0
(c) RURAL_D2
@@
We present the odometry error in Tab. I and depict the trajectory results in Fig. 6. Unexpectedly, we observe that our errors are lower than those of ORB-SLAM3 across nearly all scenarios and are approaching the level of LiDAR, or even surpassing it in some scenarios.
In static environments (U_A0, U_B0), our EFEAR-4D method achieves accurate pose estimation due to precise ego-velocity estimation, outperforming other methods. In sequence U_B0, dense vegetation challenges LiDAR, degrading LEGO-LOAM’s performance. Other radar methods show significant errors due to lack of feature extraction. In high-speed environments, rapid motion causes large gaps between the consecutive point cloud, reducing features for matching. Our method maintains low error rates due to accurate ego-velocity estimation. LEGO-LOAM’s larger FOV results in better stability and performance than radar and camera methods, such as ORB-SLAM3, which suffer from limited FOV. In dynamic environments, our method minimizes the impact of dynamic objects by removing them and leveraging weak reflections, outperforming vision-based and LiDAR-based methods. It also outperforms GICP and APDGICP by extracting features and filtering unstable points. In snowy environments, particularly in R_D2, radar-based methods outperform others as LEGO-LOAM occasionally fails to register frames, demonstrating radar’s efficacy in all weather conditions.
Fig. 7 describes the relative error of ORB-SLAM3, LEGO-LOAM, baseline radar odometry, and our radar odometry EFEAR-4D using different path lengths and speeds with different sequences, following the popular KITTI odometry evaluation protocol [31]. Our EFEAR-4D shows a stable performance on all the datasets except in high-speed scenarios where it is slightly inferior to LEGO-LOAM. Both 4DRadarSLAM (apdgicp) and ORB-SLAM3 have a relatively large error and their performances are sensitive to speed variations. It is worth pointing out that both our EFEAR-4D and 4DRadarSLAM perform well in dynamic environments, proving that 4D radar odometry is not less sensitive to moving objects than its counterparts of vision and LiDAR.